The following data is given for the Bahia Company:

Budgeted production (at 100% of normal capacity) 1,031 units
Actual production   963 units
Materials:
    Standard price per pound $1.85
    Standard pounds per completed unit 11
    Actual pounds purchased and used in production 10,275
    Actual price paid for materials $21,064
Labor:
    Standard hourly labor rate $14.08 per hour
    Standard hours allowed per completed unit 4.2
    Actual labor hours worked 4,959.45
    Actual total labor costs $75,632
Overhead:
    Actual and budgeted fixed overhead $1,080,000
    Standard variable overhead rate $28.00 per standard labor hour
    Actual variable overhead costs $138,865
Overhead is applied on standard labor hours.

Round your final answer to the nearest dollar. Do not round interim calculations.

The fixed factory overhead volume variance is

a.$71,232 unfavorable

b.$71,232 favorable

c.$25,616 unfavorable

d.$25,616 favorable
Answer #1
Fixed Factory overhead volume variance = budgeted fixed overhead-absorbed overhead
absorption rate = actual & budgeted fixed overhead/Budgeted production
absorption rate = 1080000/1031 = 1047.527
Fixed factory overhead volume variance = 1080000-(963*1047.527)
fixed factory overhead volume variance = 1080000-1008768
Fixed factory overhead volume variance = 71232 Favorable
Option "b" IS Correct

Give a thumbs up if it is helpful

please let me know if any doubt